Timothy Carter went out to eat with his girlfriend at a fancy restaurant. When he tried to pay the bill with his Mastercard credit card, he was told that the restaurant accepted only cash or American Express. His waiter suggested that he use the ATM across the street to withdraw cash using his credit card. Tim did as suggested and didn't pay attention to any fees until he received his credit card statement one month later. He was shocked to see the total fees (3.5% cash advance), and his APR was increased to 20.0%. Given the cost of the meal ($150) plus the associated fees, how much did his meal cost him?
Answer #1

APR charged = 20%

monthly APR = 0.20 / 12 months = 0.0167 monthly

cash advance = 3.5 %

cost of the meal = $ 150

calculation of interest fees = $ 150*0.0167 = $ 2.505

cash advance fees =$ 150 * 0.035 = $ 5.25

total cost of meal = 150 + 2.505 + 5.25 = $ 157.76